# Freezing of Gait Detection Using Gramian Angular Fields and Federated Learning from Wearable Sensors

### v1

@@ Abstract.
Freezing of gait (FOG) is a debilitating symptom of Parkinson’s disease (PD) that impairs mobility and safety. Traditional detection methods face challenges due to intra and inter-patient variability, and most systems are tested in controlled settings, limiting their real-world applicability. Addressing these gaps, we present FOGSense, a novel FOG detection system designed for uncontrolled, free-living conditions. It uses Gramian Angular Field (GAF) transformations and federated deep learning to capture temporal and spatial gait patterns missed by traditional methods. We evaluated our FOGSense system using a public PD dataset, ‘tdcsfog’. FOGSense improves accuracy by 10.4% over a single-axis accelerometer, reduces failure points compared to multi-sensor systems, and demonstrates robustness to missing values. The federated architecture allows personalized model adaptation and efficient smartphone synchronization during off-peak hours, making it effective for long-term monitoring as symptoms evolve. Overall, FOGSense achieves a 22.2% improvement in F1-score compared to state-of-the-art methods, along with enhanced sensitivity for FOG episode detection. Code is available: https://github.com/shovito66/FOGSense.
Wearable, Sensors, Federated Learning, CNN, Parkinson, Gait
ccs: Computer systems organization Sensor networksccs: Computing methodologies Neural networksccs: Computing methodologies Distributed algorithmsccs: Applied computing Health informatics
1. Introduction
@@
Approximately 90,000 people over the age of 65 are diagnosed with Parkinson’s disease (PD) in the U.S. every year (Willis et al., 2022). Freezing of Gait (FOG) is one of the most incapacitating symptoms of PD, a neurodegenerative condition that impairs coordination and mobility. The transient inability to start or continue walking, a hallmark of FOG, significantly impairs a patient’s mobility and quality of life, increasing the risk of falls, accidents, and major health concerns (Brederecke, 2023; Lichter et al., 2021). Therefore, timely detection of FOG and effective interventions can save a PD patient from severe injuries or even death. Timely treatments, including external cues or therapy changes, can help reduce falls, increase mobility, and improve the overall quality of life when FOG is detected early (Ahlrichs et al., 2016). Additionally, real-time detection and ongoing monitoring could keep a doctor well-informed and help modify treatment regimens, which can lessen the frequency and intensity of FOG episodes (Borzì et al., 2021).
The integration of wearable technologies with machine learning (ML) represents a major advancement in the continuous monitoring and management of PD symptoms. This convergence enables real-time motion pattern recognition, predictive modeling of FOG episodes, and personalized therapeutic insights. Conventional time series frameworks, however, struggle to capture the complex spatiotemporal dynamics and variability in patient movement, limiting their effectiveness in processing the multivariate data streams required for real-time FOG prediction and intervention (Pardoel et al., 2019; Wang and Oates, 2015).
Wearable FOG detection systems also face challenges in sensor reliability and robustness. Sensor failures in mobile health monitoring can decrease detection accuracy by up to 65% when using conventional imputation methods (Mamun et al., 2022). This vulnerability is especially critical in systems relying on multiple sensors or data axes, where each dependency increases the risk of system failure in real-world deployments (Mamun et al., 2022).
The Gramian Angular Field (GAF) transformation is a signal processing technique for FOG detection, commonly used with accelerometer and gyroscope data. By converting temporal sequences into 2D representations, GAF improves FOG detection with Convolutional Neural Networks (CNN) (Elmir et al., 2023). It outperforms wavelet and recurrence plot analysis by preserving local and global patterns while maintaining order (Wang and Oates, 2015). Mapping data to polar coordinates in the Gramian matrix, GAF captures locomotor perturbations, and its 2D CNN approach extracts richer features than 1D CNNs, boosting FOG detection accuracy (Le et al., 2023; Mamani and del Alamo, 2019).
@@
To further investigate and enhance the efficacy of GAF and CNN for FOG detection, we designed FOGSense, a novel framework for FOG episode detection using only one axis of a single accelerometer sensor. This minimalist approach reduces hardware dependencies and enhances reliability while maintaining strong detection capabilities. By applying GAF transformations with a CNN, FOGSense enables both window-level and episode-level FOG detection. It effectively captures essential gait signatures within brief time windows, allowing for real-time use on resource-limited devices with low computational demand. Additionally, its single-axis design and dynamic weight transfer method improve precision and resilience against sensor failures.
@@
Subsequently, we take another step to improve FOGSense by integrating federated learning. Modern smartphones can train models on-device but often strain battery and CPU (Ignatov et al., 2019; Yang et al., 2019). FOGSense overcomes this by employing a distributed learning framework tailored for wearable-smartphone ecosystems (Wang et al., 2020; Khan et al., 2020). In this setup, sensor data is transferred during off-peak hours to capture FOG patterns without interrupting regular use (Mazilu et al., 2015). A central server then uses TensorFlow to train models, converts them to TensorFlow Lite for deployment, and adapts these models to both individual and population-wide FOG variations (Mathur et al., 2017; Luo et al., 2020; Li et al., 2020; Rieke et al., 2020). This approach is optimized for resource-constrained devices (Mathur et al., 2017), with asynchronous updates that balance performance and efficiency (Yang et al., 2019).
@@
In summary, our contributions are (i) a review of FOG detection methods and challenges, (ii) a FOG detection solution using GAF, CNN, and federated learning with a single accelerometer, (iii) the development and evaluation of the FOGSense system for window-level, episode-level, and federated FOG detection, (iv) performance comparison across channels and with existing methods, and (v) a dynamic weight transfer framework to address channel failure.
2. Related Work
@@
Raw sensor-based deep learning minimizes manual feature engineering by applying models like 1D-CNNs directly to accelerometer data (Hannink et al., 2016; Zhao et al., 2018). These methods reduce preprocessing overhead but require large labeled datasets and often fail to capture the full complexity of sensor signals. Moreover, they struggle with generalizability, particularly when faced with variations in patient populations or sensor placements (Sama et al., 2012), limiting their real-world applicability.
SSL techniques aim to overcome labeled data constraints by learning meaningful representations from raw data (Soumma et al., 2024; Xia et al., 2024). These methods extract essential features without extensive annotation but face challenges in adapting to diverse sample distributions. Variations in environmental conditions, sensor noise, and individual gait patterns can significantly affect performance (Lonini et al., 2018), highlighting the need for robust adaptation mechanisms.
Federated learning enables collaborative model training across devices while preserving data privacy (Li et al., 2020; Chen et al., 2020). This distributed approach is particularly suitable for healthcare but presents challenges in resource-constrained wearables, such as balancing computational efficiency, power consumption, and model performance.
@@
Figure 1. Proposed distributed federated system design.
3. FOGSense System Design
3.1. Federated Learning on Edge Devices
Developing a reliable FOG detection model is challenging due to the complexity of PD and symptom variability across patients. Our hierarchical system architecture, illustrated in Fig. 1, addresses these challenges while accommodating the constraints of wearable devices. Wearables continuously collect acceleration data, which is transferred to paired smartphones at designated intervals. A central federated server aggregates this data during off-peak hours, applies GAF transformations, and executes model training. The processed data are stored in a time-series database for comprehensive health monitoring and temporal analysis. Updated model weights are saved in an in-memory database for efficient synchronization with smartphones, which then update wearables. This design ensures efficient data handling, robust model distribution, and reliable FOG monitoring.
Figure 2. System architecture illustrating the federated learning workflow: local models are trained and uploaded by devices, then aggregated into a global model, which is distributed back to each device.
3.2. Dataset
We used the publicly available PD dataset  (Howard et al., 2023) from the TLVMC FOG prediction competition. The dataset has 62 subjects diagnosed with Parkinson’s disease, specifically, focusing on FOG episodes. The analysis concentrated on the ‘tdcsfog’ dataset. Data were collected via a lower-back triaxial accelerometer at 128 Hz, capturing vertical (AccV), mediolateral (AccML), and anteroposterior (AccAP) movements, critical for identifying directional mobility loss during FOG events. Rich temporal annotations, including start hesitations, turns, and walking, enabled precise FOG analysis. Using DHWT the FOG window counts rose from 1,096,265 to 1,994,705, yielding an 82% increase.
@@ 3.2. Dataset
Figure 3. Class distribution normal and FOG events before (a) and after (b) applying Differential Hopping Windowing Technique (DHWT).
3.3. Pre-processing
The dataset, originally collected at 128Hz, was downsampled to 64Hz to reduce computational complexity while preserving signal characteristics (Soumma et al., 2024). The ‘tdcsfog’ data’s three event types were consolidated into a single column, labeling FOG as 1 and no-FOG as 0.
A critical preprocessing step involved defining time windows for analysis, which we set at 4-second intervals to effectively capture essential gait features for FOG detection (Naghavi and Wade, 2019). To address class imbalance effectively, we incorporated the Differential Hopping Windowing Technique (DHWT)(Soumma et al., 2024). This technique applied varying overlap sizes: no overlap (0%) for no-FOG windows and an increased 50% overlap for FOG windows. By using this method, we enhanced the representation of FOG events, mitigating class imbalance and enriching the model’s training dataset without additional preprocessing. Fig 3 illustrates the results achieved with this approach, showing a more balanced distribution of FOG and no-FOG instances. We filtered out impure windows where no-FOG events (class 0) predominated, designating windows where FOG events (class 1) were in the majority as pure FOG windows. For improved data consistency and distribution, we subtracted the mean value from each window ($X_{i,j}^{\prime}=X_{i,j}-\mu_{X_{i}}$), ensuring normalization across samples.
@@ 3.3. Pre-processing
3.4. Gramian Angular Fields
The GAF transformation effectively converts 1D time series data into 2D images suitable for CNN-based analysis while preserving temporal and dynamic structures (Yokkampon et al., 2022). The transformation process consists of two key steps: normalization and polar coordinate mapping. First, the time series data is normalized to the range $[-1,1]$ through the following transformations:
(1) $x^{\prime}_{i}$ $=\frac{x_{i}-\min(X)}{\max(X)-\min(X)}$ $\theta_{i}$ $=\cos^{-1}(x^{\prime}_{i})$
(2) $x^{\prime\prime}_{i}$ $=2x^{\prime}_{i}-1$ $r_{i}$ $=\frac{t_{i}}{N}$
@@ 3.4. Gramian Angular Fields
Figure 4. GAF transformation of accelerometer signals.
@@
(3) $\text{GASF}(i,j)=\cos(\theta_{i}+\theta_{j})$
@@
(4) $\text{GAF}=\begin{bmatrix}\cos(\theta_{1}+\theta_{1})&\cdots&\cos(\theta_{1}+\theta_{m})\\ \vdots&\ddots&\vdots\\ \cos(\theta_{m}+\theta_{1})&\cdots&\cos(\theta_{m}+\theta_{m})\end{bmatrix}$
@@
Figure 5. GAF-transformed images for (a) No-FOG, (b) FOG conditions, and (c) differences, with significantly different regions highlighted with red boxes.
As shown in Fig. 5, the GAF transformation creates distinct visual patterns for normal (Fig. 5a) and FOG (Fig. 5b) conditions, enabling image-based deep learning applications. The annotated difference, based on pixel-wise absolute differences between GAF-transformed images, highlights areas with significant pattern variations, capturing subtle gait changes. The top 15 regions with the highest difference values are emphasized, showcasing where the FOG pattern deviates most from no-FOG. This approach effectively captures temporal patterns in sequential data, making it ideal for medical signal processing and gait analysis (Buz et al., 2020).
3.5. Dataset Stratification
We used subject-based stratified partitioning to prevent data leakage and ensure robust evaluation (Saeb et al., 2017). The dataset was split into training (70%), validation (10%), and testing (20%) sets using randomly shuffled subject identifiers and it preserved FOG/no-FOG proportions and temporal integrity (Hammerla et al., 2016). This procedure is repeated for each group within the dataset, conducted three times, and the results are averaged to ensure reliability and robustness in our findings.
3.6. Experimental Setup
Our FOG detection system employs a sophisticated multichannel 2D Convolutional Neural Network (CNN) architecture designed to maximize performance through spatial relationship analysis across multiple sensor modalities. The architecture processes data from three distinct accelerometer axes—Vertical (AccV), Mediolateral (AccML), and Anteroposterior (AccAP)—by transforming them into GAF images and processing them through dedicated CNN branches. This approach enables the system to leverage complementary information from different sensor perspectives, significantly enhancing its FOG detection capabilities.
Figure 6. The multichannel CNN architecture.
The proposed architecture, as illustrated in Fig. 6, consists of three parallel CNN branches, each dedicated to processing GAF images from a specific accelerometer axis. Each branch processes input data with dimensions (64, 64, 3) through three consecutive convolutional blocks of increasing complexity.
Table 1. Dense layers after feature concatenation.

### v3

@@ Abstract
Freezing of gait (FOG) is a debilitating symptom of Parkinson’s disease that impairs mobility and safety by increasing the risk of falls. An effective FOG detection system must be accurate, real-time, and deployable in free-living environments to enable timely interventions. However, existing detection methods face challenges due to (1) intra- and inter-patient variability, (2) subject-specific training, (3) using multiple sensors in FOG dominant locations (e.g., ankles) leading to high failure points, (4) centralized, non-adaptive learning frameworks that sacrifice patient privacy and prevent collaborative model refinement across populations and disease progression, and (5) most systems are tested in controlled settings, limiting their real-world applicability for continuous in-home monitoring. Addressing these gaps, we present FOGSense, a real-world deployable FOG detection system designed for uncontrolled, free-living conditions using only a single sensor. FOGSense uses Gramian Angular Field (GAF) transformations and privacy-preserving federated deep learning to capture temporal and spatial gait patterns missed by traditional methods with a low false positive rate. We evaluated our system using a public Parkinson’s dataset collected in a free-living environment. FOGSense improves accuracy by 10.4% over a single-axis accelerometer, reduces failure points compared to multi-sensor systems, and demonstrates robustness to missing values. The federated architecture allows personalized model adaptation and efficient smartphone synchronization during off-peak hours, making it effective for long-term monitoring as symptoms evolve. Overall, FOGSense achieved a 22.2% improvement in F1-score and a 74.53% reduction in false positive rate compared to state-of-the-art methods, along with enhanced sensitivity for FOG episode detection, empowering preventive care and long-term symptom management as Parkinson’s progresses. Code is available: https://github.com/shovito66/FOGSense.
Index Terms— Parkinson’s Disease, Federated Learning, Wearable Sensors, Freezing of Gait, Movement Disorder, Data Scarcity, Class Imbalance, IMU
I Introduction
@@
Freezing of Gait (FOG) is one of the most incapacitating symptoms of Parkinson’s disease (PD), a neurodegenerative condition that impairs coordination and mobility. The transient inability to start or continue walking, a hallmark of FOG, significantly impairs a patient’s mobility and quality of life, increasing the risk of falls, accidents, and major health concerns [1, 2]. Therefore, timely detection of FOG and effective interventions can save a PD patient from severe injuries or even death. Timely treatments, including external cues or therapy changes, can help reduce falls, increase mobility, and improve the overall quality of life when FOG is detected early [3]. Additionally, real-time detection and ongoing monitoring could keep a doctor well-informed and help modify treatment regimens, which can lessen the frequency and intensity of FOG episodes [4].
The integration of wearables with machine learning (ML) represents a major advancement in the continuous monitoring and management of PD symptoms. This convergence enables real-time motion pattern recognition, predictive modeling of FOG episodes, and personalized therapeutic insights. However, the effectiveness of ML-driven FOG detection depends heavily on sensor placement, reliability, and adaptability, as capturing nuanced gait dynamics remains a challenge [5]. While CyclePro [6] demonstrates the utility of wearable sensors for continuous gait cycle detection, especially in noisy environments, it also highlights the necessity for platform-independent solutions that do not require predefined parameters, which could greatly benefit FOG detection systems. SmartSock [7] offers a robust, wearable platform for continuous monitoring, using machine learning and signal processing techniques to provide real-time, context-aware assessments of ankle edema.
Multi-sensor systems, while providing richer data, introduce critical reliability concerns. Wearable FOG detection solutions that rely on multiple sensors at FoG-dominant locations (e.g., ankles [8] or shoes [9]) are susceptible to sensor failures, misalignment, and increased system complexity. Sensor failures can decrease detection accuracy by up to 65% [10], especially in systems that rely on multiple channels, where each dependency increases the risk of system failure in real-world deployments.
Despite the advantages of wearables, FOG detection remains challenging due to the complex, non-linear nature of gait disturbances [11]. Conventional time series frameworks struggle to capture the complex spatiotemporal dynamics and variability in gait movement, limiting their effectiveness in processing the multivariate data streams required for real-time FOG detection and intervention [12, 13]. To address this, we leverage Gramian Angular Field (GAF) transformations, a signal processing technique that converts temporal sequences into 2D representations and improves FOG detection [14]. It outperforms wavelet and recurrence plot analysis by preserving local and global patterns while maintaining order [15], boosting FOG detection accuracy [16].
@@
To further investigate and enhance the efficacy of GAF and Convolutional Neural Network (CNN) for FOG detection, we designed FOGSense, a novel framework for FOG episode detection using only one axis of a single accelerometer sensor with a low false positive rate. This minimalist approach reduces hardware dependencies and enhances reliability while maintaining strong detection capabilities. By applying GAF transformations with a CNN, FOGSense enables both window-level and episode-level FOG detection. It effectively captures essential gait signatures within brief time windows, allowing for real-time use on resource-limited devices with low computational demand. Additionally, its single-axis design and dynamic weight transfer method improve precision and resilience against sensor failures.
@@
Subsequently, we take another step to improve FOGSense by integrating federated learning. Modern smartphones can train models on-device but often strain battery and CPU [17, 18]. FOGSense overcomes this by employing a distributed learning framework tailored for wearable-smartphone ecosystems [19, 20]. This setup transfers sensor data during off-peak hours to capture FOG patterns without interrupting regular use [21]. A central server then uses TensorFlow to train models, converts them to TensorFlow Lite for deployment, and adapts these models to both individual and population-wide FOG variations [22, 23, 24]. To achieve this, we explore the research question: How can we reduce false positive (FP) intervention rates while maintaining high FOG detection accuracy, ensuring the model adapts to different patients over time without compromising privacy?
@@
In summary, our contributions are (i) a real-world deployable, privacy-preserving FOG detection system using a single lower-back accelerometer, reducing multi-sensor failure risks, (ii) Integration of GAF transformations and federated learning, enhancing gait pattern recognition with lower FP rate (iii) the development and evaluation of the FOGSense system for window-level, episode-level, and federated FOG detection, (iv) performance comparison across channels and with existing methods, and (v) a dynamic weight transfer framework to address channel failure.
II Related Work
@@
Raw sensor-based deep learning minimizes manual feature engineering by applying models like 1D-CNNs directly to accelerometer data [25]. These methods reduce preprocessing overhead but require large labeled datasets and often fail to capture the full complexity of sensor signals. Moreover, they struggle with generalizability, particularly when faced with variations in patient populations or sensor placements [26], limiting their real-world applicability. Recent works like DGAD  [27], FoG-Finder [28] and Gait-Guard [8] use two ankle-worn IMUs and CNN-based models for real-time FoG detection. DGAD applies a Butterworth bandpass filter (0.5-8Hz) to IMU data before using a single-headed CNN for classification. While some of these systems achieve strong detection performance, they rely on more than one sensor, increasing hardware complexity and failure risks, and are tested in a controlled setting.
SSL techniques aim to overcome labeled data constraints by learning meaningful representations from raw data [29, 30]. These methods extract essential features without extensive annotation but face challenges in adapting to diverse sample distributions. Variations in environmental conditions, sensor noise, and individual gait patterns can significantly affect performance [31], highlighting the need for robust adaptation mechanisms.
Federated learning enables collaborative model training across devices while preserving data privacy [23, 32]. This distributed approach is particularly suitable for healthcare but presents challenges in resource-constrained wearables, such as balancing computational efficiency, power consumption, and model performance.
@@
III FOGSense System Design
III-A Federated Learning on Edge Devices
Developing a reliable FOG detection model is challenging due to the complexity of PD and symptom variability across patients. Our hierarchical system architecture, illustrated in Fig. 1, addresses these challenges while accommodating the constraints of wearable devices. Wearables continuously collect acceleration data, which is transferred to paired smartphones at designated intervals. A central federated server aggregates this data during off-peak hours, applies GAF transformations, and executes model training as shown in Fig. 2. The processed data are stored in a time-series database for comprehensive health monitoring and temporal analysis. Updated model weights are saved in an in-memory database for efficient synchronization with smartphones, which then update wearables. This design ensures efficient data handling, robust model distribution, and reliable FOG monitoring.
Figure 1: Proposed distributed federated system design.
We trained a multi-channel CNN model across five federated clients, each simulating different data distributions. Clients were assigned different combinations of gait data, allowing the system to learn personalized gait variations without centralizing sensitive patient data. We used a weighted averaging strategy (FedAvg) to aggregate client updates [33], ensuring a balanced contribution to the global model while preserving model stability. Each client trained on local epochs before synchronizing with the central server, which converted the models to TensorFlow Lite for efficient deployment on resource-constrained devices. This approach enables continuous, personalized adaptation of FOG detection while maintaining low communication overhead and strong privacy guarantees.
Figure 2: System architecture illustrating the federated learning workflow: local models are trained and uploaded by devices, then aggregated into a global model, which is distributed back to each device.
III-B Dataset
We used the publicly available PD dataset  [34] from the TLVMC FOG prediction competition. The dataset has 62 subjects diagnosed with Parkinson’s disease, specifically focusing on FOG episodes. The analysis concentrated on the ‘tdcsfog’ dataset. Data were collected via a lower-back triaxial accelerometer at 128 Hz, capturing vertical (AccV), mediolateral (AccML), and anteroposterior (AccAP) movements, critical for identifying directional mobility loss during FOG events. Rich temporal annotations, including start hesitations, turns, and walking, enabled precise FOG analysis. Using the “Differential Hopping Windowing Technique” (DHWT) [29], the FOG window counts rose from 1,096,265 to 1,994,705, yielding an 82% increase.
@@ III-B Dataset
Figure 3: Class distribution normal and FOG events before (a) and after (b) applying (DHWT) to the training set.
III-C Dataset Stratification
We used subject-based stratified partitioning to prevent data leakage and ensure robust evaluation. The dataset was split into training (70%), validation (10%), and testing (20%) sets using randomly shuffled subject identifiers and it preserved FOG/no-FOG proportions and temporal integrity [35]. This procedure is repeated for each group within the dataset, conducted three times, and the results are averaged to ensure reliability and robustness in our findings.
III-D Pre-processing
The dataset, originally collected at 128Hz, was downsampled to 64Hz to reduce computational complexity while preserving signal characteristics [29]. The ‘tdcsfog’ data’s three event types were consolidated into a single column, labeling FOG as 1 and no-FOG as 0.
A critical preprocessing step involved defining time windows for analysis, which we set at 4-second intervals to effectively capture essential gait features for FOG detection [36]. To address class imbalance effectively, we used DHWT [29] on our training set. This technique applied varying overlap sizes: no overlap (0%) for no-FOG windows and an increased 50% overlap for FOG windows. By using this method, we enhanced the representation of FOG events, mitigating class imbalance and enriching the model’s training dataset without additional preprocessing. Fig. 3 illustrates the results achieved with this approach, showing a more balanced distribution of FOG and no-FOG instances. We filtered out impure windows where no-FOG events (class 0) predominated, designating windows where FOG events (class 1) were in the majority as pure FOG windows. For improved data consistency and distribution, we subtracted the mean value from each window ($X_{i,j}^{\prime}=X_{i,j}-\mu_{X_{i}}$), ensuring normalization across samples.
@@ III-D Pre-processing
III-E Gramian Angular Fields
The GAF transformation effectively converts 1D time series data into 2D images suitable for CNN-based analysis while preserving temporal and dynamic structures [37]. The transformation process consists of two key steps: normalization and polar coordinate mapping. First, the time series data is normalized to the range $[-1,1]$ through the following transformations:
$x^{\prime}_{i}$ $=\frac{x_{i}-\min(X)}{\max(X)-\min(X)}$ $\theta_{i}$ $=\cos^{-1}(x^{\prime}_{i})$ (1)
$x^{\prime\prime}_{i}$ $=2x^{\prime}_{i}-1$ $r_{i}$ $=\frac{t_{i}}{N}$ (2)
@@ III-E Gramian Angular Fields
Figure 4: GAF transformation of accelerometer signals.
@@
$\text{GASF}(i,j)=\cos(\theta_{i}+\theta_{j})$ (3)
@@
$\text{GAF}=\begin{bmatrix}\cos(\theta_{1}+\theta_{1})&\cdots&\cos(\theta_{1}+\theta_{m})\\ \vdots&\ddots&\vdots\\ \cos(\theta_{m}+\theta_{1})&\cdots&\cos(\theta_{m}+\theta_{m})\end{bmatrix}$ (4)
@@
Figure 5: GAF-transformed images for (a) No-FOG, (b) FOG conditions, and (c) differences, with significantly different regions highlighted with red boxes.
As shown in Fig. 5, the GAF transformation creates distinct visual patterns for normal (Fig. 5a) and FOG (Fig. 5b) conditions, enabling image-based deep learning applications. The annotated difference, based on pixel-wise absolute differences between GAF-transformed images, highlights areas with significant pattern variations, capturing subtle gait changes. The top 15 regions with the highest difference values are emphasized, showcasing where the FOG pattern deviates most from no-FOG. This approach effectively captures temporal patterns in sequential data, making it ideal for medical signal processing and gait analysis [38].
III-F Experimental Setup
FOGSense uses a multichannel 2D-CNN architecture designed to maximize performance through spatial relationship analysis across multiple sensor modalities. The model processes data from three distinct accelerometer axes (vertical, mediolateral, and anteroposterior) by transforming them into GAF images and processing them through dedicated CNN branches. This approach enables the system to leverage complementary information from different sensor perspectives, significantly enhancing its FOG detection capabilities. All experiments, including pre-processing, model training, and federated learning, were conducted using Python 3.8, TensorFlow 2.2, and one NVIDIA A100 GPU.
Figure 6: The multichannel CNN architecture.
The proposed architecture (Fig. 6) consists of three parallel CNN branches, each processing GAF images from a specific accelerometer axis. Each branch processes input data with dimensions (64, 64, 3) through three consecutive convolutional blocks with increasing complexity - starting with 32 filters (20% dropout), followed by 64 and 128 filters, incorporating L2 regularization ($\lambda=0.001$) and dropout rates of 20% and 40%.
TABLE I: Dense layers after feature concatenation.